The pressure of a gas is reduced from 1200.0 mm Hg to 850.0 mm Hg as the volume of its container is increased by moving a piston from 85.0 mL to 350.0 milliliters. What would the final temperature be if the original temperature were 90.0 °C?


1
Expert's answer
2021-05-17T04:08:51-0400

"\frac{P_1V_1}{T_1}=\frac{P_2V_2}{T_2}"


"\frac{1200×85.0}{363}=\frac{850×350}{T_2}"


"T_2= 1058.75K=785.75\u00b0C"